# Preserving Source Video Realism: High-Fidelity Face Swapping for Cinematic Quality

### v1

@@ Abstract
Video face swapping is crucial in film and entertainment production, where achieving high fidelity and temporal consistency over long and complex video sequences remains a significant challenge. Inspired by recent advances in reference-guided image editing, we explore whether rich visual attributes from source videos can be similarly leveraged to enhance both fidelity and temporal coherence in video face swapping. Building on this insight, this work presents LivingSwap, the first video reference guided face swapping model. Our approach employs keyframes as conditioning signals to inject the target identity, enabling flexible and controllable editing. By combining keyframe conditioning with video reference guidance, the model performs temporal stitching to ensure stable identity preservation and high-fidelity reconstruction across long video sequences. To address the scarcity of data for reference-guided training, we construct a paired face-swapping dataset, Face2Face, and further reverse the data pairs to ensure reliable ground-truth supervision. Extensive experiments demonstrate that our method achieves state-of-the-art results, seamlessly integrating the target identity with the source video’s expressions, lighting, and motion, while significantly reducing manual effort in production workflows. Project webpage: https://aim-uofa.github.io/LivingSwap.
@@ 1 Introduction
Video face swapping holds significant value in the film and entertainment industries. However, existing methods fall short of meeting the stringent demands of high-quality cinematic production. For instance, GAN-based methods [li2019faceshifter, deepfake, chen2020simswap, shiohara2023blendface, luo2025canonswap], which typically operate in a frame-by-frame manner (see Fig. 2a), have made considerable progress in injecting the target identity. Yet they often struggle with realism and suffer from temporal artifacts—such as flickering and jitter—especially in long sequences. Meanwhile, contemporary video diffusion models [zhao2023diffswap, han2024face, chen2024hifivfs, wang2025dynamicface], while achieving high visual quality and temporal consistency, often rely on sparse conditioning signals such as facial landmarks. This reliance makes it challenging to perfectly align the generated expressions, lighting, and subtle nuances with the source video, resulting in faces that may appear unnatural or lack lifelike vitality. Consequently, there is a critical need for a video face swapping model capable of directly leveraging the rich, detailed information from the source video’s facial region.
@@
Achieving a high degree of customization while preserving the integrity of the original content remains a fundamental challenge in generative media [yang2023object]. Methods based on DDIM inversion [ju2023direct, qi2023fatezero, geyer2023tokenflow] or Score Distillation Sampling (SDS) [poole2022dreamfusion, hertz2023delta] often struggle to strike an optimal balance between editability and fidelity. In the field of video editing, a common strategy involves combining inpainting with structural guidance such as depth or keypoints [jiang2025vace, hu2025animate, tu2025videoanydoor] (see Fig. 2b). However, such approaches inherently discard the original pixel information within the edited region, leading to a noticeable loss of fidelity in details.
@@
Recently, reference guided generation has demonstrated remarkable breakthroughs in image editing, successfully reconciling editing flexibility with high-fidelity reconstruction [labs2025flux, deng2025emerging, wu2025qwen]. This approach directly guides the model using the reference images, enabling the faithful utilization of rich visual attributes contained in the references. Nevertheless, adapting these techniques to video face swapping presents unique challenges: (1) the difficulty of injecting a stable and consistent identity condition throughout long and complex video sequences; and (2) the scarcity of paired training data for reference-guided video face swapping task.
@@
In this work, we address these challenges by introducing LivingSwap, the first video editing model for face swapping that directly references the source video’s details. To facilitate this, we decompose the challenging task of long-video face swapping into a highly controllable pipeline comprising keyframe identity injection, video reference completion, and temporal stitching. This pipeline not only enables flexible identity guidance using high-quality image swapping results, but also mitigates the accumulation of errors that typically arises in long videos.Furthermore, we construct Face2Face, the first-of-its-kind dataset specifically curated for video reference-guided face swapping. To ensure reliable ground-truth supervision, we reverse each data pair by using the generated results as inputs and the original data as the ground truth.
@@ 2 Related Work
Video face swapping. The task of video face swapping is to replace the identity in a video while preserving attributes such as pose, expression, illumination, and background. GAN-based approaches [li2019faceshifter, deepfake, chen2020simswap, shiohara2023blendface, luo2025canonswap], which typically process videos frame-by-frame, have made notable progress in injecting target identity through encoder–decoder pipelines. However, they often suffer from temporal inconsistencies—such as flickering and jitter—especially in long sequences. Recently, video diffusion models are used in video face swapping. Diffusion-based methods  [zhao2023diffswap, han2024face, chen2024hifivfs, wang2025dynamicface] demonstrate stronger generative power and achieve higher visual quality and temporal consistency. They treat face swapping as inpainting by masking the original face and regenerating it conditioned on background frames and auxiliary attributes [chen2024hifivfs, wang2025dynamicface]. This often leads to the loss of fine-grained details and introduces inconsistencies with the model’s pretrained priors, thereby degrading generation quality. In this work, we tackle these challenges by directly leveraging detailed source video references for face swapping, combined with a carefully curated dataset Face2Face and an reversing data strategy to provide high-fidelity supervision.
Diffusion-based Video Editing. With the rapid development of diffusion models, a wide range of customization and editing techniques [ruiz2023dreambooth, cao2023masactrl, yang2024lora, zhu2024unleashing, zhao2025local, zhao20253d, yang2025any] have emerged. From a methodological perspective, these approaches can be broadly categorized into inversion-based, inpainting-based, and reference-guided methods. Inversion-based methods [ju2023direct, qi2023fatezero, geyer2023tokenflow, poole2022dreamfusion, hertz2023delta] reconstruct the original video trajectory in the diffusion process to enable editing, but they often struggle to balance editability and fidelity. Inpainting-based approaches [jiang2025vace, hu2025animate, tu2025videoanydoor] edit masked regions with structural guidance such as optical flow, depth, or keypoints, achieving temporal coherence but usually at the cost of losing fine-grained details. Recently, reference-guided methods [labs2025flux, deng2025emerging, wu2025qwen, hurst2024gpt, comanici2025gemini] have shown strong potential by leveraging reference images to combine flexible editing with high-fidelity reconstruction. Nonetheless, extending this paradigm to long video sequences remains challenging due to the scarcity of paired data and the difficulty of maintaining consistent identity or attributes over time. In this work, we use reference-guided generation for face video swapping, enabling controllable identity transfer while preserving temporal coherence and visual fidelity across long sequences.
@@ 3 Preliminary: Video Generation with DiT and Rectified Flow
Recent advancements in diffusion-based video generation leverage the Diffusion Transformer (DiT) architecture combined with continuous-time training objectives such as Rectified Flow (RF) [esser2024scaling] to achieve high-quality and temporally coherent synthesis. DiT extends traditional UNet-based diffusion backbones with transformer blocks, enabling more flexible and scalable modeling of high-dimensional video data. In this framework, the model learns a continuous denoising process by predicting the velocity between a pair of latent points. Given a ground-truth sample $x_{1}$, and a standard Gaussian noise $x_{0}\sim\mathcal{N}(0,I)$, a linearly interpolated latent $x_{t}$ is constructed as:
@@ 4.1 Keyframes Identity Injection
Maintaining a stable target identity across long and dynamic video sequences remains a fundamental challenge in video face swapping. Compared to video-level approaches, image-level face swapping methods often provide stronger and more precise identity injection. Motivated by the complementary strengths of image-based face swapping and video interpolation paradigms [wang2024framer, guo2025keyframe], we introduce a keyframe-based identity injection strategy that delivers robust and consistent identity conditioning for long-sequence video generation.
@@
These keyframes are chosen at moments exhibiting significant variations in pose, expression, or illumination, ensuring that the major appearance changes across the video are well captured. Next, we perform Per-frame Edit on these keyframes using a high-quality image-level method [facefusion2025]. This process can be optionally followed by manual refinement using tools such as Adobe Photoshop, which aligns with industrial workflows requiring both high-quality results and flexible editability.
@@
serves as a temporal boundary condition that guides the diffusion model during sequence completion. Compared to per-frame processing commonly adopted in industrial pipelines, our keyframe scheme requires modifying only a small set of boundary keyframes, which preserves the visual quality of frame-by-frame editing while substantially improving efficiency.
@@ 4.2 Video Reference Completion
Inspired by the reference-guided architecture of the video editing foundation model VACE [jiang2025vace], we extend this paradigm to video face swapping to achieve high-fidelity reconstruction. As illustrated in Fig. 3 Part 2, instead of masking the facial area in the source video and depending on external pretrained encoders, we directly feed the complete source video segment
@@
To integrate identity and appearance signals, we encode each input using the VAE encoder $\mathcal{E}_{\phi}(\cdot)$ and concatenate the resulting latent tokens in a temporally aligned order [zhao2025diceptiongeneralistdiffusionmodel]:
@@
To support adaptive feature injection, we introduce an attribute encoder composed of DiT blocks, mirroring the architecture of the diffusion backbone and initialized with matching pretrained weights [zhang2023adding, jiang2025vace]. At each layer, the output of the attribute encoder is injected into the corresponding layer of the backbone via element-wise addition, enabling hierarchical and fine-grained conditioning in latent space. Formally, the injection process is defined as:
@@ 4.4 Dataset Construction
As discussed in Sec. 4.1, the Per-frame Edit process provides high-quality single-frame face-swapped results. This industrial approach, often based on GAN-based face swapping models such as Inswapper [facefusion2025], achieve strong fidelity to the source video by leveraging the full-pixel source frame as input. However, such results struggle with realism and suffer from temporal inconsistencies and degraded visual quality, including artifacts and distortions (Fig. 5).
@@
Benefiting from the prior knowledge inherited from the pretrained model and our role-reversing strategy, LivingSwap exhibits strong robustness to noisy training samples, effectively going beyond the limitations of training data quality, as further demonstrated in Sec. 5.3 and Appendix A. Finally, leveraging Face2Face, we apply the rectified flow loss (detailed in Sec. 3) to supervise LivingSwap in preserving source-video attributes with high fidelity.
@@ 5.1 Experimental Setup
Dataset. For training, we construct our dataset Face2Face using CelebV-Text [yu2023celebv] and VFHQ [xie2022vfhq]. CelebV-Text is a large-scale video–text dataset containing approximately 70,000 in-the-wild facial video clips, totaling around 279 hours of footage. VFHQ comprises over 16,000 high-resolution facial video clips, covering diverse identities and scenarios. Based on these two datasets, we employ Inswapper [facefusion2025] to generate paired face-swapping data and then reverse the Input-GroundTruth roles to construct our final training dataset, Face2Face. The construction details of dataset please refer to Sec. 4.4 and Appendix G
@@
Table 2: Quantitative comparison with SOTA methods on FF++.
@@
Benchmark. For evaluation, we adopt FaceForensics++ (FF++) [rossler2019faceforensics++], a widely used benchmark for video face manipulation analysis. However, FF++ primarily consists of interview-style or livestream videos, which do not adequately reflect the challenging conditions commonly encountered in cinematic productions. To properly assess model performance in real film-like environments, we construct a new benchmark, CineFaceBench. CineFaceBench contains 400 target–source test pairs. The curated video clips span a wide range of challenging cinematic scenarios, including long-take shots, complex lighting conditions, exaggerated expressions, heavy facial makeup, and semi-transparent occlusions. To further assess robustness under different levels of identity similarity, each video is paired with two target images—an easy and a hard case—selected based on identity similarity scores. The construction details of CineFaceBench please refer to Appendix H.
@@
Metrics. To comprehensively evaluate the face-swapping performance, we employ both image-level and video-level metrics to assess the quality of the generated results. Following prior work [chen2020simswap, chen2024hifivfs, wang2025dynamicface], we randomly sample 10 frames from each face-swapped video to compute image-level evaluation metrics, including ID Similarity, Expression Error, Lighting Error, Gaze Error, and Face Pose Error. ID Similarity is measured by encoding both the face-swapped result and the target image into identity vectors using a pre-trained ID encoder [wang2018cosface], followed by computing the cosine similarity between them. In addition to identity-related metrics, we calculate Expression and Lighting Errors by extracting their respective coefficients using a 3DMM-based face reconstruction method [deng2019accurate] and computing the L2 distance between the source and swapped results. Similarly, we use a gaze estimation model [abdelrahman2023l2cs] with cosine similarity and a head pose estimation model [ruiz2018fine] with L2 distance to quantify the gaze and head pose changes. For video-level evaluation, we use Frechet Video Distance (FVD) [unterthiner2018towards] to assess the overall quality of generated videos. Additionally, we compute the average ranking of each method across all metrics to provide a comprehensive assessment of model performance.
@@
Implementation Details. We initialize LivingSwap with the 14B pretrained weights of VACE [jiang2025vace]. Specifically, we train the model for 10,000 steps using the AdamW optimizer, with a learning rate of 1e-5 and a batch size of 16. The input resolution is set to 640, and the number of frames is set to 81, following the original VACE configuration. The final model is trained on 8 NVIDIA H200 GPUs for 14 days. During inference, we first detect faces using a face detection model, then crop the detected regions and perform face swapping on the cropped sequences. The swapped regions are subsequently pasted back into their original positions within the frames. As a face-swapping model commonly used in industrial scenarios, we employ Inswapper [facefusion2025] as Per-frame Edit method for processing keyframes. In the ablation study, all models are trained for 2000 steps with the same hyperparameters. Moreover, we select the 100 most challenging cases in FF++ for evaluation, ensuring a fair comparison and a clearer demonstration of each ablated model’s effectiveness.

### v2

@@ Abstract
Video face swapping is crucial in film and entertainment production, where achieving high fidelity and temporal consistency over long and complex video sequences remains a significant challenge. Inspired by recent advances in reference-guided image editing, we explore whether rich visual attributes from source videos can be similarly leveraged to enhance both fidelity and temporal coherence in video face swapping. Building on this insight, this work presents LivingSwap, the first video reference guided face swapping model. Our approach employs keyframes as conditioning signals to inject the target identity, enabling flexible and controllable editing. By combining keyframe conditioning with video reference guidance, the model performs temporal stitching to ensure stable identity preservation and high-fidelity reconstruction across long video sequences. To address the scarcity of data for reference-guided training, we construct a paired face-swapping dataset, Face2Face, and further reverse the data pairs to ensure reliable ground-truth supervision. Extensive experiments demonstrate that our method achieves state-of-the-art results, seamlessly integrating the target identity with the source video’s expressions, lighting, and motion, while significantly reducing manual effort in production workflows. Project webpage: https://aim-uofa.github.io/LivingSwap
@@ 1 Introduction
Video face swapping holds significant value in the film and entertainment industries. However, existing methods fall short of meeting the stringent demands of high-quality cinematic production. For instance, GAN-based methods [24, 7, 4, 33, 25], which typically operate in a frame-by-frame manner (see Fig. 2a), have made considerable progress in injecting the target identity. Yet they often struggle with realism and suffer from temporal artifacts—such as flickering and jitter—especially in long sequences. Meanwhile, contemporary video diffusion models [52, 14, 5, 38, 49], while achieving high visual quality and temporal consistency, often rely on intermediate representations from external encoders (e.g., facial landmarks or 3D faces), which inevitably lose information compared to raw source videos. This reliance makes it challenging to perfectly align the generated expressions, lighting, and subtle nuances with the source video, resulting in faces that may appear unnatural or lack lifelike vitality. Consequently, there is a critical need for a video face swapping model capable of directly leveraging the rich, detailed information from the source video’s facial region.
@@
Achieving a high degree of customization while preserving the integrity of the original content remains a fundamental challenge in generative media [44]. Methods based on DDIM inversion [21, 29, 12] or Score Distillation Sampling (SDS) [28, 16] often struggle to strike an optimal balance between editability and fidelity. In the field of video editing, a common strategy involves combining inpainting with structural guidance such as depth or keypoints [20, 17, 34] (see Fig. 2b). However, such approaches inherently discard the original pixel information within the edited region, leading to a noticeable loss of fidelity in details.
@@
Recently, reference guided generation has demonstrated remarkable breakthroughs in image editing, successfully reconciling editing flexibility with high-fidelity reconstruction [22, 8, 40]. This approach directly guides the model using the reference images, enabling the faithful utilization of rich visual attributes contained in the references. Nevertheless, adapting these techniques to video face swapping presents unique challenges: (1) the difficulty of injecting a stable and consistent identity condition throughout long and complex video sequences; and (2) the scarcity of paired training data for reference-guided video face swapping task.
@@
In this work, we address these challenges by introducing LivingSwap, the first video editing model for face swapping that directly references the source video’s details. To facilitate this, we decompose the challenging task of long-video face swapping into a highly controllable pipeline comprising keyframe identity injection, video reference completion, and temporal stitching. This pipeline not only enables flexible identity guidance using high-quality image swapping results, but also mitigates the accumulation of errors that typically arise in long videos.Furthermore, we construct Face2Face, the first-of-its-kind dataset specifically curated for video reference-guided face swapping. To ensure reliable ground-truth supervision, we reverse each data pair by using the generated results as inputs and the original data as the ground truth.
@@ 2 Related Work
Video face swapping. The task of video face swapping is to replace the identity in a video while preserving attributes such as pose, expression, illumination, and background. GAN-based approaches [24, 7, 4, 33, 25, 35, 19], which typically process videos frame-by-frame, have made notable progress in injecting target identity through encoder–decoder pipelines. However, they often suffer from temporal inconsistencies—such as flickering and jitter—especially in long sequences. Recently, video diffusion models are used in video face swapping. Diffusion-based methods  [52, 14, 5, 38] demonstrate stronger generative power and achieve higher visual quality and temporal consistency. They treat face swapping as inpainting by masking the original face and regenerating it conditioned on background frames and auxiliary attributes [5, 38]. This often leads to the loss of fine-grained details and introduces inconsistencies with the model’s pretrained priors, thereby degrading generation quality. In this work, we tackle these challenges by directly leveraging detailed source video references for face swapping, combined with a carefully curated dataset Face2Face and a reversing data strategy to provide high-fidelity supervision.
Diffusion-based Video Editing. With the rapid development of diffusion models, a wide range of customization and editing techniques [32, 3, 42, 54, 53, 51, 43] have emerged. From a methodological perspective, these approaches can be broadly categorized into inversion-based, inpainting-based, and reference-guided methods. Inversion-based methods [21, 29, 12, 28, 16] reconstruct the original video trajectory in the diffusion process to enable editing, but they often struggle to balance editability and fidelity. Inpainting-based approaches [20, 17, 34] edit masked regions with structural guidance such as optical flow, depth, or keypoints, achieving temporal coherence but usually at the cost of losing fine-grained details. Recently, reference-guided methods [22, 8, 40, 18, 6] have shown strong potential by leveraging reference images to combine flexible editing with high-fidelity reconstruction. Nonetheless, extending this paradigm to long video sequences remains challenging due to the scarcity of paired data and the difficulty of maintaining consistent identity or attributes over time. In this work, we use reference-guided generation for face video swapping, enabling controllable identity transfer while preserving temporal coherence and visual fidelity across long sequences.
@@ 3 Preliminary: Video Generation with DiT and Rectified Flow
Recent advancements in diffusion-based video generation leverage the Diffusion Transformer (DiT) architecture combined with continuous-time training objectives such as Rectified Flow (RF) [10] to achieve high-quality and temporally coherent synthesis. DiT extends traditional UNet-based diffusion backbones with transformer blocks, enabling more flexible and scalable modeling of high-dimensional video data. In this framework, the model learns a continuous denoising process by predicting the velocity between a pair of latent points. Given a ground-truth sample $x_{1}$, and a standard Gaussian noise $x_{0}\sim\mathcal{N}(0,I)$, a linearly interpolated latent $x_{t}$ is constructed as:
@@ 4.1 Keyframes Identity Injection
Maintaining a stable target identity across long and dynamic video sequences remains a fundamental challenge in video face swapping. Compared to video-level approaches, image-level face swapping methods often provide stronger and more precise identity injection. Motivated by the complementary strengths of image-based face swapping and video interpolation paradigms [39, 13], we introduce a keyframe-based identity injection strategy that delivers robust and consistent identity conditioning for long-sequence video generation.
@@
These keyframes are chosen at moments exhibiting significant variations in pose, expression, or illumination, ensuring that the major appearance changes across the video are well captured. Next, we perform Per-frame Edit on these keyframes using a high-quality image-level method [15]. This process can be optionally followed by manual refinement using tools such as Adobe Photoshop, which aligns with industrial workflows requiring both high-quality results and flexible editability.
@@
as a temporal boundary condition that guides the diffusion model during sequence completion. Compared to per-frame processing commonly adopted in industrial pipelines, our keyframe guidance scheme requires modifying only a small set of boundary keyframes, which preserves the visual quality and flexibility of frame-by-frame editing while substantially improving efficiency.
@@ 4.2 Video Reference Completion
Inspired by the reference-guided architecture of the video editing foundation model VACE [20], we extend this paradigm to video face swapping to achieve high-fidelity reconstruction. As illustrated in Fig. 3 Part 2, instead of masking the facial area in the source video and depending on external pretrained encoders, we directly feed the complete source video segment
@@
To integrate identity and appearance signals, we encode each input using the VAE encoder $\mathcal{E}_{\phi}(\cdot)$ and concatenate the resulting latent tokens in a temporally aligned order [50]:
@@
To support adaptive feature injection, we introduce an attribute encoder composed of DiT blocks, mirroring the architecture of the diffusion backbone and initialized with matching pretrained weights [48, 20]. At each layer, the output of the attribute encoder is injected into the corresponding layer of the backbone via element-wise addition, enabling hierarchical and fine-grained conditioning in latent space. Formally, the injection process is defined as:
@@ 4.4 Dataset Construction
As discussed in Sec. 4.1, the Per-frame Edit process provides high-quality single-frame face-swapped results. This industrial approach, often based on GAN-based face swapping models such as Inswapper [15], achieves strong fidelity to the source video by leveraging the full-pixel source frame as input. However, such results struggle with realism and suffer from temporal inconsistencies and degraded visual quality, including artifacts and distortions (Fig. 5).
@@
Benefiting from the prior knowledge inherited from the pretrained model and our role-reversing strategy, LivingSwap exhibits strong robustness to noisy training samples, effectively going beyond the limitations of training data quality, as further demonstrated in Sec. 5.3. Finally, leveraging Face2Face, we apply the rectified flow loss (detailed in Sec. 3) to supervise LivingSwap in preserving source-video attributes with high fidelity.
@@ 5.1 Experimental Setup
Dataset. For training, we construct our dataset Face2Face using CelebV-Text [47] and VFHQ [41]. CelebV-Text is a large-scale video–text dataset containing approximately 70,000 in-the-wild facial video clips, totaling around 279 hours of footage. VFHQ comprises over 16,000 high-resolution facial video clips, covering diverse identities and scenarios. Based on these two datasets, we employ Inswapper [15] to generate paired face-swapping data and then reverse the Input-GroundTruth roles to construct our final training dataset, Face2Face. Please refer to Sec. 4.4 and the Supplementary Material for construction details.
@@
Table 2: Quantitative comparison with state-of-the-art methods on FF++.
@@
Benchmark. For evaluation, we adopt FaceForensics++ (FF++) [30], a widely used benchmark for video face manipulation analysis. However, FF++ primarily consists of interview-style or livestream videos, which do not adequately reflect the challenging conditions commonly encountered in cinematic productions. To properly assess model performance in real film-like environments, we construct a new benchmark, CineFaceBench. CineFaceBench contains 400 target–source test pairs. The curated video clips span a wide range of challenging cinematic scenarios, including long-take shots, complex lighting conditions, exaggerated expressions, heavy facial makeup, and semi-transparent occlusions. To further assess robustness under different levels of identity similarity, each video is paired with two target images—an easy and a hard case—selected based on identity similarity scores. Please refer to the Supplementary Material for construction details.
@@
Metrics. To comprehensively evaluate the face-swapping performance, we employ both image-level and video-level metrics to assess the quality of the generated results. Following prior work [4, 5, 38], we randomly sample 10 frames from each face-swapped video to compute image-level evaluation metrics, including ID Similarity, Expression Error, Lighting Error, Gaze Error, and Face Pose Error. ID Similarity is measured by encoding both the face-swapped result and the target image into identity vectors using a pre-trained ID encoder [37], followed by computing the cosine similarity between them. In addition to identity-related metrics, we calculate Expression and Lighting Errors by extracting their respective coefficients using a 3DMM-based face reconstruction method [9] and computing the L2 distance between the source and swapped results. Similarly, we use a gaze estimation model [1] with cosine similarity and a head pose estimation model [31] with L2 distance to quantify the gaze and head pose changes. For video-level evaluation, we use Frechet Video Distance (FVD) [36] to assess the overall quality of generated videos. Additionally, we compute the average ranking of each method across all metrics to provide a comprehensive assessment of model performance.
@@
Implementation Details. We initialize LivingSwap with the 14B pretrained weights from [20]. The model is trained for 10,000 steps using the AdamW optimizer with a learning rate of 1e-5 and a batch size of 16. Following the original VACE configuration, the input resolution is set to 640 and the number of frames is set to 81. The final model is trained on 8 NVIDIA H200 GPUs for approximately 14 days. During inference, we first detect faces using a pre-trained face detector, crop the detected regions, and perform face swapping on the cropped sequences. The swapped regions are then pasted back into their original locations in the frames. To ensure reproducibility and fair comparison, we adopt a fixed-interval keyframe policy (every 79 frames for 81-frame chunks) in all experiments. As a face-swapping model widely used in industrial applications, we employ Inswapper [15] as the per-frame editing method for processing keyframes, consistent with the Face2Face construction. For the ablation study, all models are trained for 2,000 steps with the same hyperparameters. In addition, we select the 100 most challenging samples from FF++ for evaluation to better highlight the effectiveness of each ablated variant.